A 2.0-kg block oscillates on the end of a mass le
#22. Answer is A.
Please show steps, thank you!
Answer #1

Time Period from the graph is T= 4.0s.

Mass of the block, m= 2.0kg

Now if "k" is the spring constant then Time Period is given by

  T= 2 \pi \sqrt{\frac{m}{k}}

Using all the given values in above,

4.0= 2 \times 3.14 \sqrt{\frac{2.0}{k}}

\sqrt{\frac{2.0}{k}}= 0.637

Squaring both sides, we get

\frac{2.0}{k}= 0.406

or Spring constant, k= \frac{2.0}{0.406}= 4.926 N/m

Now at time t= 1.0s, displacement is x= 0.080m

then Potential Energy of the system at this time will be

P.E= \frac{1}{2}kx^2= \frac{1}{2} \times 4.926 \times (0.08)^2

  P.E= 0.01576 J\sim 0.016J(ANS)